Polyester has a Young's, Modulus of 3.5 GPa and PLGA 1.2 GPa. A polymer blend is prepared with 45:55 mixture (PET:PlGA) . What is the Young's, Modulus of this blend (use mixture rule)



1
Expert's answer
2022-04-13T19:36:04-0400

Youngs modulus = YPET = 3.5 GPa, YGPA = 1.2 GPa

It has mixture of 45:50

Ym = YPET * 0.45 + YGPA * 0.55 = 3.5 * 0.45 + 1.2 * 0.55 = 2.24 GPa